OpenStudy (anonymous):

NO ONE CANT ANSWER THIS ! Scott currently has an account balance of $2,147.39. He opened the account five years ago with a deposit of $1,852.10. If the interest compounds monthly, what is the interest rate on the account?

OpenStudy (whpalmer4):

\[F = P(1+\frac{i}{n})^{nt}\]where \(P\) is present value, \(F\) is future value, \(i\) is annual nominal interest rate, expressed as a decimal, \(n\) is the number of compounding periods per year, \(t\) is the number of years.

OpenStudy (anonymous):

2147.39=1852.10*(1+x)^60 (since compound monthly, you have exponent (5*12) 2147.39/1852.10=(1+x)^60 (divide 1852.10 on both sides) 1.1594=(1+x)^60, add ln (natural log)on both sides ln1.1594=ln(1+x)^60, 0.1479=60ln(1+x), divide 60 on both sides, 0.002465=ln(1+x), so e^(0.002465)=1+x, x=0.002468, so monthly interest is 0.002468, annual interest=0.002468*12=0.03=3%